If the consumption function is reading as C=290+0.9Y. Therefore the savings function will read as

1
Expert's answer
2021-12-08T10:38:24-0500

S=f(Y)

Y=C+S

S=Y-C

Thus savings is the amount of income that is not spent on consumption.

C=a+bY……….i

S=Y-a-bY………ii

=-a+Y(1-b)

Where b= MPC

           a= Autonomous consumption

Using the consumption function given,

C=290+0.9Y

From Equation ii above,

S=Y-290-0.9Y

Therefore, S= -290+0.1Y